# Verification and Refinement of Natural Language Explanations through LLM-Symbolic Theorem Proving

### v1

@@ 3 Explanation-Refiner
To verify the logical validity and refine any logical errors in explanatory sentences for NLI tasks, we present a neuro-symbolic framework, Explanation-Refiner, to iteratively check and refine the explanation $E_{i}$ based on external feedback; Isabelle (Paulson, 1994) and its proof assistant Isabelle/HOL (Nipkow et al., 2002) are applied as external tools to make the logical deduction and provide feedback for explanation refinement.
Figure 1 shows an overview of our proposed framework. Given an NLI task, to evaluate the logical validity of the entailment, the LLM is prompted to perform an autoformalisation process that transforms natural language sentences into formal language represented in the form of Isabelle theory. Each fact $f\in E_{i}$ is converted into an axiom $a_{i}$, where each $a_{i}$ is an element of the set $A=\{a_{1},a_{2},...,a_{n}\}$. The premise $p_{i}$ and corresponding hypothesis $h_{i}$, is converted into a theorem for proving $p_{i}\wedge B\rightarrow h_{i}$, where $B\subseteq A$. A syntax refinement mechanism is subsequently applied to the previously transferred symbolic forms. The theorem prover is implemented as a checker to identify any syntax errors and provide these error details as feedback to an LLM, enabling the LLM to iteratively correct the syntax errors over a fixed number of iterations, denoted by $t$.
@@
Figure 2: An example of representing the premise and hypothesis sentences in Isabelle theorem
@@
Figure 3: An example of proof constructed by the Isabelle/HOL proof assistant to verify the hypothesis.
@@ 3.1 Autoformalisation
Isabelle Theory Construction
For the Isabelle theorem prover, a theory script is essential to facilitate the proof of a theorem. Therefore, we designate explanatory sentences as axioms:
@@ 3.2 Proof Construction
A proof provides a detailed, step-by-step strategy that elucidates the logical connections and unification among axioms to support the reasoning process aimed at achieving the solver’s goal. Initially, we prompt the LLM to create a preliminary proof to assess how it infers the hypothesis and to identify which explanatory sentences are relevant, redundant, or unrelated. Based on this initial proof, we then guide the LLM to develop an Isabelle proof (figure 3) that utilise the Isabelle/HOL proof assistant to clearly demonstrate the explanatory sentences (axioms) required to prove the hypothesis. Details of proof construction process prompts are described in Appendix A.4.2.
@@ 3.3 Verify and Refine
Finally, the constructed theory, which includes axioms, theorems, and proof steps, is submitted to the theorem prover for verification. If the theory is validated, it outputs a logically sound explanation. If the proof fails or timeouts, we extract the first error from the solver’s error message, identify the corresponding proof step, and locate the related explanatory sentences (axioms) from the theory. We begin by removing redundant and irrelevant facts that are not present in the preceding Isabelle proof steps or are declared as such in the text inference strategy. Then, we prompt the LLM to refine the explanatory sentences by providing it with the error message, the failed proof step, the associated proof strategy, and the relevant explanatory sentences for further iteration. This process is iterative and progressive; with each iteration, the framework addresses one or more logical errors, continually refining the explanatory sentences to ultimately yield a logically valid and verifiable explanation. Additional details on the prompts used for refinement are described in Appendix A.4.3.
@@ 4.2 Theorem Prover
We adopted Isabelle (Paulson, 1994) as our theorem prover and Isabelle/HOL (Nipkow et al., 2002) as the proof assistant. To integrate this theorem prover as a real-time verification tool with LLMs, we employ a Python client (Shminke, 2022) as TCP (Transmission Control Protocol) client to configure Isabelle as a server. This enables the communication of the constructed theory files and the extraction of the response messages from Isabelle.
@@ A.2 Scalability
Figure 10 shows the average Isabelle solving time against the number of planned explanatory sentences in a proof and the length of suggested proof steps, including theories that have syntax errors, respectively. In some cases, the theorem prover may get stuck on a proof step, and we have set a termination time if the solving time exceeds 65 seconds.
Figure 10: (a) Average Isabelle solving time against number of explanatory sentences planned in a proof. (b) Average Isabelle solving time against number of suggested proof steps in a proof.
@@ A.4.1 Autoformalisation
Figure 14 displays the prompts used to identify action verbs (events) within the premise, explanation, and hypothesis sentences, representing events in Davidsonian-event semantics. Figure 15 displays the prompts used to transfer natural language to logical forms based on the identified events verbs. Figure 16 shows how to convert logical forms into Isabelle code (axioms and type declaration). Figure 17 shows how to convert the premise and hypothesis sentences into the Isabelle theorem code, based on the previously constructed axioms code. Figure 18 shows how to refine the syntax errors based on the types of errors, the provided code, the error messages, and the locations of the errors within the code.
@@ A.5 Examples of Explanation Refinement
Table 1 shows an example from the e-SNLI dataset of how the explanation changes after each iteration. Figures 22, 23, and 24 illustrate the Isabelle theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.2 QASC Refined Examples
Table 4 shows an example from the QASC dataset of how the explanation changes after each iteration. Figures 30, 31 illustrate the Isabelle theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.3 WorldTree Refined Examples
Table 7 shows an example from the WorldTree dataset of how the explanation changes after each iteration. Figures 36, 37, 38, 39, 40, 41, 42 and 43 illustrate the Isabelle theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.4 Unrefined Example
Table 10 and Table 11 shows an example from the WorldTree dataset that does not refine within 10 iterations and is not caused by a syntax error. The figures that follow show the detailed Isabelle theory code of the related iterations.
@@
Input : Premise $p$, Explanation $E$, Hypothesis $h$, Isabelle server $isabelle$, Autoformalisation model $m_{a}$, Isabelle syntax refinement model $m_{sr}$, Rough inference model $m_{ri}$, Proof step build model $m_{pr}$, Facts filter model $m_{f}$, Explanation refinement model $m_{e}$
@@ 44      $iterations\leftarrow iterations+1$
Figure 16: Prompts for converting logical form into Isabelle code format for building the axioms and type declaration
Figure 17: Prompts for building the theorem code part of the Isabelle theory
@@
Figure 20: Prompts for how to build a proof for Isabelle proof assistant

### v2

@@ 3 Explanation-Refiner
To verify the logical validity and refine any logical errors in explanatory sentences for NLI tasks, we present a neuro-symbolic framework, Explanation-Refiner, to iteratively check and refine the explanation $E_{i}$ based on external feedback; Isabelle/HOL (Nipkow et al., 2002) proof assistant is applied as the external tool to make the logical deduction and provide feedback for explanation refinement.
Figure 1 shows an overview of our proposed framework. Given an NLI task, to evaluate the logical validity of the entailment, the LLM is prompted to perform an autoformalisation process that transforms natural language sentences into formal language represented in the form of Isabelle/HOL theory. Each fact $f\in E_{i}$ is converted into an axiom $a_{i}$, where each $a_{i}$ is an element of the set $A=\{a_{1},a_{2},...,a_{n}\}$. The premise $p_{i}$ and corresponding hypothesis $h_{i}$, is converted into a theorem for proving $p_{i}\wedge B\rightarrow h_{i}$, where $B\subseteq A$. A syntax refinement mechanism is subsequently applied to the previously transferred symbolic forms. The theorem prover is implemented as a checker to identify any syntax errors and provide these error details as feedback to an LLM, enabling the LLM to iteratively correct the syntax errors over a fixed number of iterations, denoted by $t$.
@@
Figure 2: An example of representing the premise and hypothesis sentences in Isabelle/HOL theorem
@@
Figure 3: An example of proof constructed by the LLM for the Isabelle/HOL proof assistant to verify the hypothesis.
@@ 3.1 Autoformalisation
Isabelle/HOL Theory Construction
For the Isabelle/HOL theorem prover, a theory script is essential to facilitate the proof of a theorem. Therefore, we designate explanatory sentences as axioms:
@@ 3.2 Proof Construction
A proof provides a detailed, step-by-step strategy that elucidates the logical connections and unification among axioms to support the reasoning process aimed at achieving the solver’s goal. Initially, we prompt the LLM to create a preliminary proof in natural language to assess how it infers the hypothesis and to identify which explanatory sentences are relevant, redundant, or unrelated. Based on this initial proof, we then guide the LLM to develop an Isabelle/HOL proof (figure 3) that utilise the Isabelle/HOL proof assistant to clearly demonstrate the explanatory sentences (axioms) required to prove the hypothesis. The general proof steps generated by an LLM are in the format ‘show $X$ using $Y$ by $Z$’, where the theorem prover is asked to prove $X$ given the assumptions $Y$, using the automated proof tactic $Z$. The proof tactic often applied is ‘blast’, which is one of Isabelle’s powerful FOL theorem proving tactics, enabling efficient and automated proof discovery across a range of logical forms (Paulson, 1999). Details of proof construction process prompts are described in Appendix A.4.2.
@@ 3.3 Verify and Refine
Finally, the constructed theory, which includes axioms, theorems, and proof steps, is submitted to the theorem prover for verification. If the theory is validated, it outputs a logically sound explanation. If the proof fails or timeouts, we extract the first error from the solver’s error message, identify the corresponding proof step, and locate the related explanatory sentences (axioms) from the theory. We begin by removing redundant and irrelevant facts that are not present in the preceding Isabelle/HOL proof steps or are declared as such in the text inference strategy. Then, we prompt the LLM to refine the explanatory sentences by providing it with the error message, the failed proof step, the associated proof strategy, and the relevant explanatory sentences for further iteration. This process is iterative and progressive; with each iteration, the framework addresses one or more logical errors, continually refining the explanatory sentences to ultimately yield a logically valid and verifiable explanation. Additional details on the prompts used for refinement are described in Appendix A.4.3.
@@ 4.2 Theorem Prover
We adopted Isabelle/HOL (Nipkow et al., 2002) proof assistant as the theorem prover. To integrate this theorem prover as a real-time verification tool with LLMs, we employ a Python client (Shminke, 2022) as TCP (Transmission Control Protocol) client to configure Isabelle/HOL as a server. This enables the communication of the constructed theory files and the extraction of the response messages from Isabelle.
@@ A.2 Scalability
Figure 10 shows the average Isabelle/HOL solving time against the number of planned explanatory sentences in a proof and the length of suggested proof steps, including theories that have syntax errors, respectively. In some cases, the theorem prover may get stuck on a proof step, and we have set a termination time if the solving time exceeds 65 seconds.
Figure 10: (a) Average Isabelle/HOL solving time against number of explanatory sentences planned in a proof. (b) Average Isabelle/HOL solving time against number of suggested proof steps in a proof.
@@ A.4.1 Autoformalisation
Figure 14 displays the prompts used to identify action verbs (events) within the premise, explanation, and hypothesis sentences, representing events in Davidsonian-event semantics. Figure 15 displays the prompts used to transfer natural language to logical forms based on the identified events verbs. Figure 16 shows how to convert logical forms into Isabelle/HOL code (axioms and type declaration). Figure 17 shows how to convert the premise and hypothesis sentences into the Isabelle/HOL theorem code, based on the previously constructed axioms code. Figure 18 shows how to refine the syntax errors based on the types of errors, the provided code, the error messages, and the locations of the errors within the code.
@@ A.5 Examples of Explanation Refinement
Table 1 shows an example from the e-SNLI dataset of how the explanation changes after each iteration. Figures 22, 23, and 24 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.2 QASC Refined Examples
Table 4 shows an example from the QASC dataset of how the explanation changes after each iteration. Figures 30, 31 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.3 WorldTree Refined Examples
Table 7 shows an example from the WorldTree dataset of how the explanation changes after each iteration. Figures 36, 37, 38, 39, 40, 41, 42 and 43 illustrate the Isabelle/HOL theory code changes during the refinement process. Green code indicates the proof steps that have successfully progressed, while red code shows where the proof failed at that step.
@@ A.5.4 Unrefined Example
Table 10 and Table 11 shows an example from the WorldTree dataset that does not refine within 10 iterations and is not caused by a syntax error. The figures that follow show the detailed Isabelle/HOL theory code of the related iterations.
@@
Input : Premise $p$, Explanation $E$, Hypothesis $h$, Isabelle//HOL server $isabelle$, Autoformalisation model $m_{a}$, Isabelle syntax refinement model $m_{sr}$, Rough inference model $m_{ri}$, Proof step build model $m_{pr}$, Facts filter model $m_{f}$, Explanation refinement model $m_{e}$
@@ 44      $iterations\leftarrow iterations+1$
Figure 16: Prompts for converting logical form into Isabelle/HOL code format for building the axioms and type declaration
Figure 17: Prompts for building the theorem code part of the Isabelle/HOL theory
@@
Figure 20: Prompts for how to build a proof for Isabelle/HOL proof assistant